OpenStudy (babyslapmafro):

Please help me solve the given differential equation by separation of variables. (4y+yx^2)dy=(2x+xy^2)dx

OpenStudy (babyslapmafro):

\[(4y+yx^2)dy=(2x+xy^2)dx\]

OpenStudy (anonymous):

ok this is a fun one :)

OpenStudy (anonymous):

first you want to pull out the variables from each parenthesis, \[(4y + yx ^{2})dy = y(4 + x ^{2}) dy \] and \[(2x + xy ^{2}) dx = x(2 + y ^{2})\]

OpenStudy (anonymous):

Forgot the dx on 2nd part

OpenStudy (anonymous):

Than the expression becomes \[y(4 + x^{2}) dy = x(2+y^{2}) dx\] divide so that X's are on right and Y's are on left

OpenStudy (anonymous):

This gives you: \[\left( \frac{ y }{ 2 + y^{2} } \right) dy = \left( \frac{ x }{ 4+x^{2} } \right) dx\]

OpenStudy (babyslapmafro):

oh ok thank you, I believe I can go from here. should all solutions of separation of variable problems be set equal to y?

OpenStudy (anonymous):

Well you want to get the y's with the dy and the x's with the dx

OpenStudy (anonymous):

but most cases you are solving for a function that is of the form y(x)

OpenStudy (anonymous):

y(x) = expression

OpenStudy (babyslapmafro):

right, but I'm talking about the final solution

OpenStudy (babyslapmafro):

right ok